OpenStudy (anonymous):

how to factor 6x^2-9x-10

OpenStudy (anonymous):

Step by step solution : Step 1 : Simplify 6x2-9x - 10

OpenStudy (imstuck):

Either the quadratic formula or complete the square.

OpenStudy (anonymous):

Solve Quadratic Equation by Completing The Square 2.2 Solving 6x2-9x-10 = 0 by Completing The Square . Divide both sides of the equation by 6 to have 1 as the coefficient of the first term : x2-(3/2)x-(5/3) = 0 Add 5/3 to both side of the equation : x2-(3/2)x = 5/3 Now the clever bit: Take the coefficient of x , which is 3/2 , divide by two, giving 3/4 , and finally square it giving 9/16 Add 9/16 to both sides of the equation : On the right hand side we have : 5/3 + 9/16 The common denominator of the two fractions is 48 Adding (80/48)+(27/48) gives 107/48 So adding to both sides we finally get : x2-(3/2)x+(9/16) = 107/48 Adding 9/16 has completed the left hand side into a perfect square : x2-(3/2)x+(9/16) = (x-(3/4)) • (x-(3/4)) = (x-(3/4))2 Things which are equal to the same thing are also equal to one another. Since x2-(3/2)x+(9/16) = 107/48 and x2-(3/2)x+(9/16) = (x-(3/4))2 then, according to the law of transitivity, (x-(3/4))2 = 107/48 We'll refer to this Equation as Eq. #2.2.1 The Square Root Principle says that When two things are equal, their square roots are equal. Note that the square root of (x-(3/4))2 is (x-(3/4))2/2 = (x-(3/4))1 = x-(3/4) Now, applying the Square Root Principle to Eq. #2.2.1 we get: x-(3/4) = √ 107/48 Add 3/4 to both sides to obtain: x = 3/4 + √ 107/48 Since a square root has two values, one positive and the other negative x2 - (3/2)x - (5/3) = 0 has two solutions: x = 3/4 + √ 107/48 or x = 3/4 - √ 107/48 Note that √ 107/48 can be written as √ 107 / √ 48 Solve Quadratic Equation using the Quadratic Formula 2.3 Solving 6x2-9x-10 = 0 by the Quadratic Formula . According to the Quadratic Formula, x , the solution for Ax2+Bx+C = 0 , where A, B and C are numbers, often called coefficients, is given by : - B ± √ B2-4AC x = ———————— 2A In our case, A = 6 B = -9 C = -10 Accordingly, B2 - 4AC = 81 - (-240) = 321 Applying the quadratic formula : 9 ± √ 321 x = ————— 12 √ 321 , rounded to 4 decimal digits, is 17.9165 So now we are looking at: x = ( 9 ± 17.916 ) / 12 Two real solutions: x =(9+√321)/12=3/4+1/12√ 321 = 2.243 or: x =(9-√321)/12=3/4-1/12√ 321 = -0.743 Two solutions were found : x =(9-√321)/12=3/4-1/12√ 321 = -0.743 x =(9+√321)/12=3/4+1/12√ 321 = 2.243